### v1

# Key polynomials in terms of ultrametric balls

Abstract.
In this paper we present characterizations of the sets of key polynomials and abstract key polynomials for a valuation $\mu$ of $K(x)$, in terms of (ultrametric) balls in the algebraic closure $\overline{K}$ of $K$ with respect to $v$, a fixed extension of $\mu_{\mid K}$ to $\overline{K}$. In particular, we show that the ways of augmenting of $\mu$, in the sense of Mac Lane, are in one-to-one correspondence with the partition of a fixed closed ball $B(a,\delta)$ associated to $\mu$ into the disjoint union of open balls $B^{\circ}(a_{i},\delta)$, modulo the action of the decomposition group of $v$. We also present a similar characterization for the set of limit key polynomials for an increasing family of valuations of $K(x)$.
Key words and phrases:
ultrametric balls, key polynomials, abstract key polynomials, limit key polynomials
2020 Mathematics Subject Classification:
Primary 13A18
Partially supported by grant PID2020-116542GB-I00 funded by the Spanish MCIN/AEI. During the realization of this project the second author was supported by a grant from Fundação de Amparo à Pesquisa do Estado de São Paulo (process numbers 2017/17835-9) and a grant from Conselho Nacional de Desenvolvimento Científico e Tecnológico (process number 303215/2022-4).
1. Introduction
For a valued field $(K,v)$ we consider a fixed extension of $v$ (which we denote again by $v$) to the algebraic closure $\overline{K}$ of $K$. Set $\Gamma=v\overline{K}$, which is the divisible hull of $vK$, and consider a fixed embedding $\Gamma\hookrightarrow\Lambda$ into a divisible ordered abelian group $\Lambda$. For $a\in\overline{K}$ and $\delta\in\Lambda$, we define the closed and open balls with center on $a$ and radius $\delta$ by
$B(a,\delta)=\{b\in\overline{K}\mid v(b-a)\geq\delta\}\,\supseteq\,B^{\circ}(a,\delta)=\{b\in\overline{K}\mid v(b-a)>\delta\},$
respectively. We consider the set of such balls in $\overline{K}$:
$\mathbb{B}:=\{B(a,\delta)\mid(a,\delta)\in\overline{K}\times\Lambda\},\;\;\mathbb{B}^{\circ}:=\{B^{\circ}(a,\delta)\mid(a,\delta)\in\overline{K}\times\Lambda\}.$
Let $\mathcal{V}=\mathcal{V}(v,\Lambda)$ be the set of all extensions of $v$ to $K(x)$, taking values in $\Lambda$. We will denote by ${\rm KP}(\mu)$ the set of key polynomials for a valuation-transcendental $\mu\in\mathcal{V}$ (Definition 2.1). We also denote by $\Psi(\mu)$ the set of abstract key polynomials for an arbitrary $\mu\in\mathcal{V}$ (Definition 2.7). Finally, ${\rm KP}_{\infty}(\mathcal{C})$ will be the set of limit key polynomials for an increasing family $\mathcal{C}$ of valuations in $\mathcal{V}$ (Definition 2.11).
The main goal of this paper is to describe the sets ${\rm KP}(\mu)$, $\Psi(\mu)$ and ${\rm KP}_{\infty}(\mathcal{C})$ in terms of elements of $\mathbb{B}$ and $\mathbb{B}^{\circ}$.
The motivation for this work comes from the following reasoning. If $K$ is algebraically closed, then both sets ${\rm KP}(\mu)$ and $\Psi(\mu)$ are subsets of $\{x-a\mid a\in K\}$ (because every polynomial in these sets is irreducible and monic). Also, in this case for every valuation-transcendental valuation $\mu\in\mathcal{V}$ there exists $(a,\delta)\in K\times\Lambda$ such that $\mu=v_{a,\delta}$ where
(1) $v_{a,\delta}\left(a_{0}+a_{1}(x-a)+\ldots+a_{r}(x-a)^{r}\right):=\min_{0\leq k\leq r}\{v(a_{k})+k\delta\}.$
In particular, for algebraically closed valued fields, the situation is much easier to handle. This was the leitmotif of some classical papers (for instance, [3, 4, 5, 6, 12, 25]), where some properties of valuations on a polynomial ring $K[x]$ were deduced from an analysis of their extensions to $\overline{K}[x]$.
This paper provides a precise and more complete picture of this approach, in what concerns the description of the sets ${\rm KP}(\mu)$, $\Psi(\mu)$ and ${\rm KP}_{\infty}(\mathcal{C})$.
The notion of key polynomial was first introduced by Mac Lane in [13] in order to describe all the extensions of a rank one discrete valuation $v$ on a field $K$ to the field of rational functions $K(x)$. The main idea of these objects is the following. For a given valuation $\mu\in\mathcal{V}$ there might exist $\phi\in K[x]$ and $\eta\in\mathcal{V}$ such that
(2) $\mu(\phi)<\eta(\phi)\mbox{ and }\mu(f)\leq\eta(f)\mbox{ for every }f\in K[x].$
If that happens we write $\mu<\eta$. Key polynomials for $\mu$ are polynomials $\phi$ of minimal degree satisfying property (2) (for a fixed $\eta$). Such polynomials allow us to construct ordinary augmentations of $\mu$ (roughly speaking, these augmentations are valuations $\eta$ that are “minimal” among valuations satisfying (2)). By doing this, one can construct a “tree” of valuations as follows: let $\mathcal{V}_{0}$ be the set of all valuations on $K(x)$ which are of the form (1). Each valuation $\mu_{0}\in\mathcal{V}_{0}$ can be augmented by using elements of ${\rm KP}(\mu_{0})$. Denote by $\mathcal{V}_{1}$ the set of all valuations obtained on this way. We can iterate this process and construct similar sets $\mathcal{V}_{n}$ for every $n\in\mathbb{N}$.
We can also consider stable limit valuations. Namely, consider an infinite family of ordinary augmentations
$\mu_{0}\longrightarrow\mu_{1}\longrightarrow\ldots\longrightarrow\mu_{n}\longrightarrow$
and suppose that for every $f\in K[x]$ there exists $n_{f}\in\mathbb{N}$ such that
$\mu_{n}(f)=\mu_{n_{f}}(f)\mbox{ for every }n\geq n_{f}.$
Then the map $\mu(f):=\mu_{n_{f}}(f)$ also defines a valuation on $K(x)$. Denote the set of all these valuations by $\mathcal{V}_{\infty}$.
The main goal of Mac Lane’s work was to answer the following question:
Question 1.1.
Is it true that
$\mathcal{V}=\left(\bigcup_{n\in\mathbb{N}_{0}}\mathcal{V}_{n}\right)\cup\mathcal{V}_{\infty}?$
Mac Lane showed that if $vK$ is discrete of rank one, then the answer to the above question is affirmative [13, Theorem 8.1].
In order to prove a similar result for the case when $vK$ is not discrete or has larger rank, Vaquié introduced the notion of limit key polynomial. With this notion, one can construct limit augmentations, which are the equivalent of ordinary augmentation for increasing family of valuations that do not have a stable limit. Vaquié showed in [26] that if one interacts the constructions using both, limit and ordinary augmentations, then one can obtain the whole set $\mathcal{V}$.
Abstract key polynomials were introduced in [8] and [21] (in [21] they are called key polynomials). They are related to the ideas presented in [9] and [10]. These objects are closely related to those presented by Mac Lane and Vaquié. The main idea is that, instead of building a set of valuations (for instance, the sets $\mathcal{V}_{n}$, $n\in\mathbb{N}_{0}\cup\{\infty\}$ as before) and checking whether all the valuations in $\mathcal{V}$ are of such form, abstract key polynomials allow us to determine the relevant polynomials in the construction of a fixed element in $\mathcal{V}$.
The above discussion illustrates the difference between the sets ${\rm KP}(\mu)$ and $\Psi(\mu)$. While the polynomials in ${\rm KP}(\mu)$ allow us to find another valuation $\eta$ for which $\mu\leq\eta$, the elements in $\Psi(\mu)$ allow us to determine valuations $\nu$ such that $\nu\leq\mu$.
For a given valuation-transcendental $\mu\in\mathcal{V}$ we can fix an extension $\overline{\mu}$ of $\mu$ to $\overline{K}(x)$. Then $\overline{\mu}$ is of the form (1) for some $a\in\overline{K}$, $\delta\in\Lambda$. The idea of using elements of $\mathbb{B}$ to describe valuations comes from the fact that if $(a,\delta),(b,\epsilon)\in\overline{K}\times\Lambda$, then
$v_{a,\delta}(f)\leq v_{b,\epsilon}(f)\mbox{ for all }f\in\overline{K}[x]\ \Longleftrightarrow\ \delta\leq\epsilon\ \mbox{ and }\ B(a,\delta)\supseteq B(b,\epsilon)\ .$
Suppose that $\mu=(v_{a,\delta})_{\mid K[x]}$ for some $a\in\overline{K}$, $\delta\in\Lambda$. The valuations $\eta$ obtained from key polynomials for $\mu$ have the property $\mu\leq\eta$. Hence, it is natural to try to obtain ${\rm KP}(\mu)$ from elements in $\mathbb{B}$ or $\mathbb{B}^{\circ}$ contained in $B(a,\delta)$. On the other hand, the valuations $\nu$ obtained from abstract key polynomials for $\mu$ satisfy $\nu\leq\mu$. Hence, it is natural to try to obtain $\Psi(\mu)$ from elements in $\mathbb{B}$ or $\mathbb{B}^{\circ}$ containing $B(a,\delta)$.
Take any such ball $B$ (containing or contained in $B(a,\delta)$). A natural candidate to be key polynomial or abstract key polynomial is the minimal polynomial of $b$ over $K$ for some $b\in B$. However, this does not have to be the right object, since there might be elements in $B$ with arbitrary degree over $K$. Hence, we look for minimal polynomials (over $K$) of elements of $B$ whose degree over $K$ is minimal.
The first main result of this paper (Theorem 4.2) tells us that $\Psi(\mu)$ is exactly the set of all minimal polynomials of elements of minimal degree over $K$, belonging to balls $B\in\mathbb{B}$ such that $B\supseteq B(a,\delta)$, where $(v_{a,\delta})_{\mid K[x]}\leq\mu$. In Section 4.2, we improve this result to obtain complete sets of abstract key polynomials for a given $\mu$, that do not have “redundant” elements.
The second set of results is presented in Section 5 and deals with key polynomials for $\mu=(v_{a,\delta})_{\mid K[x]}$. The main result (Theorem 5.6) says that for $c\in B(a,\delta)$, the minimal polynomial of $c$ belongs to ${\rm KP}(\mu)$ if and only if $c$ has minimal degree over $K$ among all the elements in $B^{\circ}(c,\delta)$. An alternative way of seeing this is the following. Let
$B(a,\delta)=\bigsqcup_{i\in I}B^{\circ}(a_{i},\delta)$
be the partition of $B(a,\delta)$ as a disjoint union of open balls of radius $\delta$. For each $i\in I$, let $\operatorname{KP}_{i}(\mu)$ be the set of minimal polynomials of the elements in $B^{\circ}(a_{i},\delta)$ of minimal degree over $K$. Then,
${\rm KP}(\mu)=\bigcup_{i\in I}{\rm KP}_{i}(\mu).$
We also present a criterion (Lemma 5.9) to determine when ${\rm KP}_{i}(\mu)={\rm KP}_{j}(\mu)$ for different $i,j\in I$. One important tool used to prove the results in Section 5 is Theorem 3.1, which can be seen as a “going-up” and “going-down” property for valuations on $K(x)$ (and their extensions to $\overline{K}(x)$).
We also deal with limit key polynomials for a family $\mathcal{C}=\{\mu_{i}\}_{i\in I}$ of increasing valuations in $\mathcal{V}$. The first relevant result (Theorem 3.14) is that for each such family, there exists another increasing family $\overline{\mathcal{C}}=\{\overline{\mu}_{i}\}_{i\in I}$ of valuations on $\overline{K}(x)$ such that $(\overline{\mu}_{i})_{\mid K[x]}=\mu_{i}$ for every $i\in I$. In particular, if each $\overline{\mu}_{i}$ is of the form $v_{a_{i},\delta_{i}}$ for some $a_{i}\in\overline{K}$, $\delta_{i}\in\Lambda$, then the balls $\left\{B(a_{i},\delta_{i})\right\}_{i\in I}$ form a descending chain, and we can easily deduce (Theorem 6.2) that ${\rm KP}_{\infty}(\mathcal{C})$ is the set of minimal polynomials over $K$ of elements of minimal degree over $K$ belonging to the ball
$B:=\bigcap_{i\in I}B(a_{i},\delta_{i}).$
Finally, we use the results in this paper to obtain a Mac Lane–Vaquié chain for a valuation in $\mathcal{V}$ (Theorem 7.2). This is done by using the construction of $\mu$-optimal ultrametric balls appearing in Section 4.2.
We emphasise that this paper was motivated by the results of [7], [18], [20] and specially [27]. In particular, some of the results in this paper already appear (at least in particular cases), or can be deduced from the results appearing there.
2. Preliminaries
Let $\mathcal{V}=\mathcal{V}(v,\Lambda)$ be the tree of all $\Lambda$-valued extensions of $v$ to the field $K(x)$. We identify every $\mu\in\mathcal{V}$ with a valuation on the polynomial ring,
$\mu\colon K[x]\longrightarrow\Lambda\cup\{\infty\},$
with trivial support; that is, $\mu^{-1}(\infty)=\{0\}$. The residue field of $(K,v)$ will be denoted by $Kv$. The value group and residue field of $\mu$ will be denoted by $\Gamma_{\mu}$ and $k_{\mu}$, respectively.
This set $\mathcal{V}$ has a partial ordering. For $\mu,\eta\in\mathcal{V}$, we define
$\mu\leq\eta\ \quad\Longleftrightarrow\quad\ \mu(f)\leq\eta(f)\quad\mbox{for all }f\in K[x].$
We say that $\mathcal{V}$ is a tree because the intervals
$(-\infty,\mu]=\left\{\nu\in\mathcal{V}\mid\nu\leq\mu\right\}$
are totally ordered for all $\mu\in\mathcal{V}$ [2].
Any node $\mu\in\mathcal{V}$ is (exclusively) of one of the following types:
  • Value-transcendental: $\Gamma_{\mu}/vK$ is not a torsion group.
  • Residue-transcendental: the extension $k_{\mu}/K\!v$ is transcendental.
  • Valuation-algebraic: $\Gamma_{\mu}/vK$ is a torsion group and $k_{\mu}/K\!v$ is algebraic.
We say that $\mu$ is valuation-transcendental if it is value-transcendental or residue-transcendental. These are precisely the inner (non-maximal) nodes of $\mathcal{V}$.
2.1. Key polynomials
The graded algebra of $\mu\in\mathcal{V}$ is the integral domain $\mathcal{G}_{\mu}=\bigoplus_{\alpha\in\Gamma_{\mu}}\mathcal{P}_{\alpha}/\mathcal{P}_{\alpha}^{+}$, where
$\mathcal{P}_{\alpha}=\{f\in K[x]\mid\mu(f)\geq\alpha\}\supseteq\mathcal{P}_{\alpha}^{+}=\{f\in K[x]\mid\mu(f)>\alpha\}.$
Every $f\in K[x]\setminus\{0\}$ has a homogeneous initial coefficient $\operatorname{in}_{\mu}f\in\mathcal{G}_{\mu}$, defined as the image of $f$ in $\mathcal{P}_{\mu(f)}/\mathcal{P}_{\mu(f)}^{+}\subseteq\mathcal{G}_{\mu}$.
Definition 2.1.
A monic $\phi\in K[x]$ is a key polynomial for $\mu$ if $(\operatorname{in}_{\mu}\phi)\mathcal{G}_{\mu}$ is a homogeneous prime ideal containing no initial coefficient  $\operatorname{in}_{\mu}f$ with $\deg f<\deg\phi$.
We denote by $\operatorname{KP}(\mu)$ the set of all key polynomials for $\mu$. These polynomials are necessarily irreducible in $K[x]$. Let us recall a criterion for the existence of key polynomials [14, Theorem 4.4], [15, Theorem 2.3].
Theorem 2.2.
For every $\mu\in\mathcal{V}$ the following conditions are equivalent.
  1. (a)
    $\operatorname{KP}(\mu)=\emptyset$.
  2. (b)
    $\mathcal{G}_{\mu}$ is a simple algebra (all nonzero homogeneous elements are units).
  3. (c)
    $\mu$ is valuation-algebraic.
  4. (d)
    $\mu$ is a leaf (maximal node) of $\mathcal{V}$.
Definition. Let $\mu\in\mathcal{V}$ be valuation-transcendental. We define the degree of $\mu$ as the smallest degree of the key polynomials for $\mu$. We denote it by $\deg(\mu)$.
The degree function preserves the ordering in $\mathcal{V}$. If $\mu<\nu$ are both valuation-transcendental, then $\deg(\mu)\leq\deg(\nu)$ [2, Lemma 2.2].
On the set $\operatorname{KP}(\mu)$ we consider the following equivalence relation:
$f\sim_{\mu}g\quad\Longleftrightarrow\quad\operatorname{in}_{\mu}f=\operatorname{in}_{\mu}g\quad\Longleftrightarrow\quad\mu(f-g)>\mu(f).$
Definition 2.3.
Let $\mu,\eta\in\mathcal{V}$, with $\mu<\eta$. We define the tangent direction of $\mu$ determined by $\eta$ as the set $\mathbf{t}(\mu,\eta)$ of all monic polynomials $\phi\in K[x]$ of smallest degree satisfying $\mu(\phi)<\eta(\phi)$.
The following basic properties of the set $\mathbf{t}(\mu,\eta)$ are proved in [26, Theorem 1.15] (see also [25, Lemma 5.3]) and [15, Corollaries 2.5 + 2.6].
Lemma 2.4.
Let $\mu$, $\eta$ be valuations in $\mathcal{V}$ such that $\mu<\eta$. Let $\phi\in\mathbf{t}(\mu,\eta)$.
  1. (i)
    The polynomial $\phi$ belongs to $\operatorname{KP}(\mu)$ and $\mathbf{t}(\mu,\eta)=\{\chi\in\operatorname{KP}(\mu)\mid\phi\sim_{\mu}\chi\}$.
  2. (ii)
    For all nonzero $f\in K[x]$, we have
    • $\mu(f)<\eta(f)\quad\Longleftrightarrow\quad\operatorname{in}_{\mu}\phi\mid\operatorname{in}_{\mu}f\ \mbox{ in }\ \mathcal{G}_{\mu}$.
    • $\mu(f)=\eta(f)\;\quad\Longrightarrow\quad\operatorname{in}_{\eta}f\mbox{ is a unit in }\ \mathcal{G}_{\eta}$.
A tangent direction is a kind of “germ of augmentations” of $\mu$. The following result gives a precise meaning for this terminology.
Lemma 2.5.
Let $\mu,\eta,\nu\in\mathcal{V}$ such that $\mu<\eta$ and $\mu<\nu$.
  1. (i)
    $\,\mathbf{t}(\mu,\eta)=\mathbf{t}(\mu,\nu)$  if and only if $\,(\mu,\eta]\cap(\mu,\nu]\neq\emptyset$.
  2. (ii)
    If $\,\mathbf{t}(\mu,\eta)\cap\mathbf{t}(\mu,\nu)\neq\emptyset$, then $\,\mathbf{t}(\mu,\eta)=\mathbf{t}(\mu,\nu)$.
Proof.
Item (i) is proven in [2, Proposition 2.4]. Item (ii) is an immediate consequence of Lemma 2.4,(i). ∎
2.2. Abstract key polynomials
For all $s\in\mathbb{N}$, the $s$-th Hasse-Schmidt derivative $\partial_{s}$ on $K[x]$ is defined by:
$f(x+y)=\sum\nolimits_{0\leq s}(\partial_{s}f)y^{s}\quad\mbox{for all }\,f\in K[x],$
where $y$ is another indeterminate.
Take $\mu\in\mathcal{V}$. If $f\in K[x]$ is non-constant, we define
$\epsilon_{\mu}(f)=\max\left\{\dfrac{\mu(f)-\mu(\partial_{s}{f})}{s}\ \Big{|}\ s\in\mathbb{N}\right\}\in\Lambda.$
Let us recall [20, Prop. 3.1], which clarifies the meaning of this value.
Theorem 2.6.
Let $\overline{\mu}$ be an arbitrary extension to $\overline{K}[x]$ of some $\mu\in\mathcal{V}$. Then, for every non-constant $f\in K[x]$ we have
$\epsilon_{\mu}(f)=\max\{\overline{\mu}(x-a)\mid a\in\operatorname{Z}(f)\},$
where $\operatorname{Z}(f)$ is the set of roots of $f$ in $\overline{K}$.
We say that $a\in\operatorname{Z}(f)$ is an optimizing root of $f$ if it satisfies $\epsilon_{\mu}(f)=\overline{\mu}(x-a)$ for some extension $\overline{\mu}$ of $\mu$ to $\overline{K}[x]$.
Definition 2.7.
A monic $Q\in K[x]$ is said to be an abstract key polynomial for $\mu$ if for every non-constant $f\in K[x]$ we have
$\deg f<\deg Q\quad\Longrightarrow\quad\epsilon_{\mu}(f)<\epsilon_{\mu}(Q).$
We denote by $\Psi(\mu)$ the set of all abstract key polynomials for $\mu$.
For every polynomial $Q\in K[x]$ consider the truncated function $\mu_{Q}$ on $K[x]$, defined as follows on $Q$-expansions:
$f=\sum\nolimits_{i\geq 0}f_{i}\,Q^{i},\ \deg f_{i}<\deg Q\ \quad\Longrightarrow\quad\ \mu_{Q}(f)=\min\{\mu(f_{i}\,Q^{i})\mid i\geq 0\}.$
For every abstract key polynomial $Q$, the truncated function $\mu_{Q}$ is a valuation such that $\mu_{Q}\leq\mu$ [21, Proposition 2.6].
Lemma 2.8.
[1, Corollary 2.22] Let $\mu\in\mathcal{V}$ and let $Q\in K[x]$ be a monic polynomial. The following conditions are equivalent.
  1. (i)
    $Q\in\Psi(\mu)$ and $\mu_{Q}=\mu$.
  2. (ii)
    $Q\in\operatorname{KP}(\mu)$ and has minimal degree in this set.
2.3. Minimal pairs
Take $(a,\delta)\in\overline{K}\times\Lambda$. The function $v_{a,\delta}$ defined in (1) is a valuation on $\overline{K}[x]$, called monomial, admitting $x-a$ as a key polynomial. For any other pair $(b,\epsilon)\in\overline{K}\times\Lambda$ we have
$v_{a,\delta}=v_{b,\epsilon}\quad\Longleftrightarrow\quad\delta=\epsilon\ \mbox{ and }\ b\in B(a,\delta).$
If $\delta$ and $\epsilon$ belong to $\Gamma$, then this condition is equivalent to $B(a,\delta)=B(b,\epsilon)$.
It is well known that every valuation-transcendental $\mu\in\mathcal{V}$ is the restriction of $v_{a,\delta}$ to $K[x]$ for some $(a,\delta)\in\overline{K}\times\Lambda$ (see for example [3, 4, 6]). Moreover, we have
$\begin{array}[]{lcl}\mu\mbox{ is residue-transcendental}&\ \Longleftrightarrow&\delta\in\Gamma,\\ \mu\mbox{ is value-transcendental}&\ \Longleftrightarrow&\delta\in\Lambda\setminus\Gamma.\end{array}$
For every $c\in\overline{K}$, let $p_{K}(c)\in K[x]$ be the minimal polynomial of $c$ over $K$. For every subset $S\subseteq\overline{K}$, we denote
$p_{K}(S)=\{p_{K}(c)\mid c\in S\}\subseteq K[x].$
Also, consider the following notation
$\begin{array}[]{rcl}\deg_{K}c&:=&\deg\left(p_{K}(c)\right)\\[5.0pt] \deg_{K}S&:=&\min\{\deg_{K}c\mid c\in S\}\\[5.0pt] \operatorname{Min}_{K}S&:=&\{c\in S\mid\deg_{K}c=\deg_{K}S\}\end{array}$
The following definition comes from [5].
Definition 2.9.
The pair $(a,\delta)\in\overline{K}\times\Lambda$ is called a minimal pair over $K$ if $a\in\operatorname{Min}_{K}B(a,\delta)$. In this case, we say that $(a,\delta)$ is a minimal pair of definition of the valuation $(v_{a,\delta})_{\mid K[x]}$.
Let us recall a characterization of abstract key polynomials in terms of minimal pairs which follows from the combination of [20, Theorem 1.1] and Lemma 2.8.
Theorem 2.10.
Let $\mu\in\mathcal{V}$. Take a monic, irreducible $Q\in K[x]$ and let $a\in\overline{K}$ be an optimizing root of $Q$. Let $\delta=\epsilon_{\mu}(Q)$. Then,
  1. (i)
    $Q\in\Psi(\mu)$ if and only if $(a,\delta)$ is a minimal pair.
  2. (ii)
    $Q\in\operatorname{KP}(\mu)$ and has minimal degree in this set if and only if $(a,\delta)$ is a minimal pair of definition of $\mu$.
2.4. Limit key polynomials
We consider now an increasing family $\mathcal{C}=\{\mu_{i}\}_{i\in I}\subset\mathcal{V}$. This means that $I$ is a totally ordered set and for every $i,j\in I$, $i<j$, we have $\mu_{i}<\mu_{j}$. A polynomial $f$ is said to be $\mathcal{C}$-stable if there exists $i_{0}\in I$ such that
$\mu_{i}(f)=\mu_{i_{0}}(f)\ \mbox{ for every }i\in I,\ i\geq i_{0}.$
Otherwise, it is called $\mathcal{C}$-unstable, in which case, we have
$\mu_{i}(f)<\mu_{j}(f)\ \mbox{ for every }i,j\in I\mbox{ with }i<j.$
Definition 2.11.
A monic polynomial $\phi$ is called a limit key polynomial for $\mathcal{C}$ if it is $\mathcal{C}$-unstable and has the smallest degree among $\mathcal{C}$-unstable polynomials. We denote by $\operatorname{KP}_{\infty}(\mathcal{C})$ the set of all limit key polynomials for $\mathcal{C}$.
If all polynomials are $\mathcal{C}$-stable, then $\operatorname{KP}_{\infty}(\mathcal{C})=\emptyset$ and $\mathcal{C}$ determines a limit valuation $\mu=\lim_{i\in I}\mu_{i}=\lim(\mathcal{C})$ as follows:
(3) $\mu(f):=\max\{\mu_{i}(f)\mid i\in I\}.$
In this case, we say that $\mu$ is the stable limit of $\mathcal{C}$, and we define
$\deg(\mu):=\max\{\deg(\mu_{i})\mid i\in I\}.$
We agree that $\deg(\mu)=\infty$ if the right-hand set is unbounded.
A valuation $\mu\in\mathcal{V}$ is the stable limit of some increasing family if and only if it is valuation-algebraic [2, Proposition 4.1].
3. Lifting chains of valuations
Consider the tree $\overline{\mathcal{V}}$ of all $\Lambda$-valued valuations on the field $\overline{K}(x)$, whose restriction to $\overline{K}$ is $v$. We have a natural restriction map:
$\operatorname{res}_{K}\colon\overline{\mathcal{V}}\longrightarrow\mathcal{V},\qquad\overline{\mu}\longmapsto\operatorname{res}_{K}(\overline{\mu}):=\overline{\mu}_{\mid K[x]}.$
Whenever we say that a valuation $\overline{\mu}$ on $\overline{K}(x)$ is an extension of some $\mu\in\mathcal{V}$, we mean that $\overline{\mu}\in\overline{\mathcal{V}}$ and $\operatorname{res}_{K}(\overline{\mu})=\mu$; that is, $\overline{\mu}$ is a common extension of $\mu$ and $v$.
The main goal of this section is to prove the following result.
Theorem 3.1.
Let $\nu<\mu\,\leq\,\eta$ be valuations on $K[x]$ and let $\overline{\mu}$ be an extension of $\mu$ to $\overline{K}[x]$. Then, there exist extensions $\overline{\nu}$ of $\nu$, and $\overline{\eta}$ of $\eta$, to $\overline{K}[x]$ such that
$\overline{\nu}<\overline{\mu}\,\leq\,\overline{\eta}.$
Theorem 3.1 can be presented in two parts: “from large to small” and “from small to large”. The “from large to small” part is easy and can be found in [19].
Lemma 3.2.
[19, Lemma 2.8] Suppose that $\nu<\mu$ in $\mathcal{V}$ and $\overline{\mu}\in\overline{\mathcal{V}}$ is an extension of $\mu$. Then, there exists an extension $\overline{\nu}\in\overline{\mathcal{V}}$ of $\nu$ such that $\overline{\nu}<\overline{\mu}$.
Thus, the real aim of this Section is to prove the “from small to large” part.
Theorem 3.3.
Suppose that $\mu<\eta$ in $\mathcal{V}$ and $\overline{\mu}\in\overline{\mathcal{V}}$ is an extension of $\mu$. Then, there exists an extension $\overline{\eta}\in\overline{\mathcal{V}}$ of $\eta$ such that $\overline{\mu}<\overline{\eta}$.
This result was proved by Vaquié in the particular situation in which $(K,v)$ is Henselian and $\eta$ is an ordinary augmentation of $\mu$ [27, Thm. 3.18].
For the proof of Theorem 3.3, we will need some preliminary results.
3.1. Ubication of roots of key polynomials
For some pair $(a,\delta)\in\overline{K}\times\Lambda$, let $\overline{\mu}=v_{a,\delta}$ and $\mu=\operatorname{res}_{K}(\overline{\mu})$. For every $c\in\overline{K}$ and $f\in K[x]$, the following equivalences were shown in [16, Lemma 2.4] and [27, Proposition 3.3], respectively:
(4) $\begin{array}[]{rcl}\operatorname{in}_{\overline{\mu}}(x-c)\ \mbox{ is a unit in }\ \mathcal{G}_{\overline{\mu}}&\Longleftrightarrow&\overline{\mu}(x-c)<\delta.\\ \operatorname{in}_{\overline{\mu}}f\ \mbox{ is a unit in }\ \mathcal{G}_{\overline{\mu}}&\Longleftrightarrow&\operatorname{in}_{\mu}f\ \mbox{ is a unitin }\ \mathcal{G}_{\mu}.\end{array}$
As mentioned in Section 2.3, every valuation-transcendental $\mu\in\mathcal{V}$ admits pairs $(a,\delta)\in\overline{K}\times\Lambda$ such that $\mu=\operatorname{res}_{K}(v_{a,\delta})$. All these pairs have a common value of $\delta$. Indeed, by Theorem 2.6, we have
$\mu=\operatorname{res}_{K}(v_{a,\delta})\quad\Longrightarrow\quad\delta=\max\left\{v_{a,\delta}(x-c)\mid c\in\overline{K}\right\}=\max\left(\epsilon_{\mu}(K[x])\right).$
This common “radius” $\delta$ will be denoted by $\delta(\mu):=\max\left(\epsilon_{\mu}(K[x])\right)$ from now on. The following result gives some information about the possible “centers” $a\in\overline{K}$.
Proposition 3.4.
Let $\mu\in\mathcal{V}$ be valuation-transcendental and denote $\delta=\delta(\mu)$. Then, for every $f\in K[x]$ the following conditions are equivalent.
  1. (a)
    $\operatorname{in}_{\mu}f$ is not a unit in $\mathcal{G}_{\mu}$.
  2. (b)
    There exists $a\in\operatorname{Z}(f)$ such that $\operatorname{res}_{K}(v_{a,\delta})=\mu$.
  3. (c)
    $\epsilon_{\mu}(f)=\delta$.
Proof.
Take any $\overline{\mu}=v_{b,\delta}$ such that $\operatorname{res}_{K}(\overline{\mu})=\mu$. Let us prove that (a) $\Rightarrow$ (b). Suppose that $\operatorname{in}_{\mu}f$ is not a unit in $\mathcal{G}_{\mu}$. By (4), $\operatorname{in}_{\overline{\mu}}f$ is not a unit in $\mathcal{G}_{\overline{\mu}}$. Hence, there exists $a\in\operatorname{Z}(f)$ such that $\operatorname{in}_{\overline{\mu}}(x-a)$ is not a unit in $\mathcal{G}_{\overline{\mu}}$; thus, $\overline{\mu}(x-a)=\delta$, again by (4). Hence, $\overline{\mu}=v_{a,\delta}$.
The implication (b) $\Rightarrow$ (c) follows immediately from Theorem 2.6.
Finally, let us show that (c) $\Rightarrow$ (a). If $\epsilon_{\mu}(f)=\delta$, then $\overline{\mu}(x-a)=\delta$ for some $a\in\operatorname{Z}(f)$, by Theorem 2.6. By (4), we deduce that $\operatorname{in}_{\overline{\mu}}(x-a)$ is not a unit in $\mathcal{G}_{\overline{\mu}}$, $\operatorname{in}_{\overline{\mu}}f$ is not a unit in $\mathcal{G}_{\overline{\mu}}$, and $\operatorname{in}_{\mu}f$ is not a unit in $\mathcal{G}_{\mu}$. ∎
Corollary 3.5.
Let $\mu=\operatorname{res}_{K}(v_{a,\delta})$ for some $(a,\delta)\in\overline{K}\times\Lambda$ and $g\in K[x]$ non-constant. Then $\epsilon_{\mu}(g)<\delta\Leftrightarrow Z(g)\cap B(a,\delta)=\emptyset$; moreover, if either one of these conditions holds, then $\mu(g)=v(g(a))$.
Roughly speaking, the corollary says that the value $v_{a,\delta}(g)$ is equal to the value of $g$ at the center $a$ of the ball $B(a,\delta)$ precisely when $g$ has no roots in $B(a,\delta)$. More precisely, the value of $g(b)$ is constant for every $b$ contained in the ball $B(a,\delta)$.
Proof.
Let $\overline{\mu}=v_{a,\delta}$. If $\epsilon_{\mu}(g)<\delta$ then for every $\alpha\in Z(g)$ we have
$v(a-\alpha)=\overline{\mu}(a-x+x-\alpha)=\overline{\mu}(x-\alpha)<\delta$
so $Z(g)\cap B(a,\delta)=\emptyset$. Conversely, if $Z(g)\cap B(a,\delta)=\emptyset$, then for every $\alpha\in Z(g)$ we have
$\overline{\mu}(x-\alpha)=\overline{\mu}(x-a+a-\alpha)=v(a-\alpha)<\delta$
For the last claim, it follows from above that we have
$\mu(g)=\sum_{b\in\operatorname{Z}(g)}\overline{\mu}(x-b)=\sum_{b\in\operatorname{Z}(g)}v(a-b)=v(g(a)).$
Corollary 3.6.
If $\mu<\eta$ in $\mathcal{V}$ are both valuation-transcendental, then $\delta(\mu)<\delta(\eta)$.
Proof.
Take any $\chi\in\mathbf{t}(\mu,\eta)$. By Proposition 3.4, $\epsilon_{\mu}(\chi)=\delta(\mu)$. Since $\mu(\chi)<\eta(\chi)$ and $\mu(f)=\eta(f)$ for all nonzero $f\in K[x]$ of smaller degree, we conclude that $\delta(\mu)=\epsilon_{\mu}(\chi)<\epsilon_{\eta}(\chi)\leq\delta(\eta)$. ∎
3.2. Rigidity
The valuation $v$ on $\overline{K}$ determines a henselization $(K^{h},v^{h})$ of $(K,v)$. In the Galois group $G={\rm Aut}(\overline{K}/K)$, consider the decomposition subgroup
$\mathcal{D}=\left\{\sigma\in G\mid v\circ\sigma=v\right\}.$
Let $K^{\operatorname{sep}}$ be the separable closure of $K$ in $\overline{K}$. The field $K^{h}\subseteq K^{\operatorname{sep}}$ is defined as the fixed field of the restriction to $K^{\operatorname{sep}}$ of the decomposition group. The valuation $v^{h}$ is just the restriction of $v$ to $K^{h}$.
Let $\mathcal{V}^{h}=\mathcal{V}(v^{h},\Lambda)$ be the tree of all $\Lambda$-valued extensions of $v^{h}$ to $K^{h}(x)$. The “rigidity theorem” [16, Thm. A] is essential for our purpose.
Theorem 3.7.
The restriction mapping $\mathcal{V}^{h}\to\mathcal{V}$, defined by $\rho\mapsto\rho_{\mid K[x]}$, is an isomorphism of posets.
The group $\mathcal{D}=\operatorname{Aut}(\overline{K}/K^{h})$ has a natural action on $\overline{K}[x]$, just by acting on the coefficients of the polynomials. Since $v\circ\sigma=v$, for all $\sigma\in\mathcal{D}$, we deduce the following action on the valuation-transcendental nodes of $\overline{\mathcal{V}}$:
(5) $v_{a,\delta}\circ\sigma^{-1}=v_{\sigma(a),\delta}\quad\mbox{for all}\quad(a,\delta)\in\overline{K}\times\Lambda,\ \sigma\in\mathcal{D}.$
Clearly, all pairs in the same $\mathcal{D}$-orbit have the same restriction to $K^{h}[x]$. We deduce from Theorem 3.7 and Proposition 3.4 the following criterion.
Proposition 3.8.
Let $a,b\in\overline{K}$ and $\delta\in\Lambda$. Then,
$\operatorname{res}_{K}(v_{a,\delta})=\operatorname{res}_{K}(v_{b,\delta})\quad\Longleftrightarrow\quad\sigma(b)\in B(a,\delta)\ \mbox{ for some }\ \sigma\in\mathcal{D}.$
Proof.
By Theorem 3.7, we can assume that $(K,v)$ is Henselian; under this assumption, $\mathcal{D}$ is precisely the group $\text{Aut}(\overline{K}/K)$.
Suppose that $\operatorname{res}_{K}(v_{a,\delta})=\operatorname{res}_{K}(v_{b,\delta})$ and let $f=p_{K}(b)$. Since $\epsilon_{\mu}(f)=\delta$ by Proposition 3.4, Theorem 2.6 shows that $v_{a,\delta}(x-\sigma(b))=\delta$, for some $\sigma\in\mathcal{D}$. Hence, $v_{a,\delta}=v_{\sigma(b),\delta}$ and so $\delta\leq v(a-\sigma(b))$.
For the opposite implication, suppose that $\delta\leq v(a-\sigma(b))$ for some $\sigma\in\mathcal{D}$. Then $v_{a,\delta}=v_{\sigma(b),\delta}=v_{b,\delta}\circ\sigma^{-1}$ (where the last equality holds by (5)). Hence, $v_{a,\delta}$ and $v_{b,\delta}$ have the same restriction to $K[x]$. ∎
Corollary 3.9.
Let $f\in\operatorname{KP}(\mu)$ for a valuation-transcendental $\mu\in\mathcal{V}$. The set of optimizing roots of $f$ is equal to $\operatorname{Z}(F)$ for some irreducible factor $F$ of $f$ in $K^{h}[x]$.
Proof.
By Proposition 3.4, $\epsilon_{\mu}(f)=\delta(\mu)$ because $\operatorname{in}_{\mu}f$ is a prime element in $\mathcal{G}_{\mu}$.
By Theorem 3.7, $\mu$ has a unique extension $\mu^{h}$ to $K^{h}[x]$. Thus, $\delta(\mu^{h})=\delta(\mu)$ is the common radius of all the extensions of $\mu$ and $\mu^{h}$ to $\overline{K}[x]$.
By [17, Proposition 5.6], there is an irreducible factor $F\in K^{h}[x]$ of $f$ such that
$\operatorname{in}_{\mu^{h}}F\in\operatorname{KP}(\mu^{h})\quad\mbox{and}\quad\operatorname{in}_{\mu^{h}}(f/F)\ \mbox{ is a unit in }\ \mathcal{G}_{\mu^{h}}.$
Hence, $\epsilon_{\mu^{h}}(f/F)<\delta(\mu^{h})=\delta(\mu)=\epsilon_{\mu}(f)$ by Proposition 3.4 applied to $\mu^{h}$. This implies that no root of $f/F$ is optimizing root of $f$. Thus, the optimizing roots of $f$ are all included in $\operatorname{Z}(F)$. Since the decomposition group acts transitively on $\operatorname{Z}(F)$, Proposition 3.8 shows that all roots of $F$ are optimizing roots of $f$. ∎
We can summarize these arguments as follows.
Remark 3.10.
Let $\mu\in\mathcal{V}$ be valuation-transcendental. There are a finite number of valuations $v_{a,\delta}$ on $\overline{K}[x]$ such that $\operatorname{res}_{K}(v_{a,\delta})=\mu$. The decomposition group $\mathcal{D}$ acts transitively on them. These valuations can be obtained from any $f\in\operatorname{KP}(\mu)$ as follows: take $\delta=\epsilon_{\mu}(f)$ and all $a\in\operatorname{Z}(F)$, where $F$ is the irreducible factor of $f$ in $K^{h}[x]$ containing all optimizing roots of $f$.
Proposition 3.8 leads to the following useful reformulation of Proposition 3.4.
Proposition 3.11.
Let $f\in K[x]$ and $\mu=\operatorname{res}_{K}(v_{a,\delta})$ for some pair $(a,\delta)\in\overline{K}\times\Lambda$. Then, $\operatorname{in}_{\mu}f$ is not a unit in $\mathcal{G}_{\mu}$ if and only if $\,\operatorname{Z}(f)\cap B(a,\delta)\neq\emptyset$.
Proof.
If $\operatorname{in}_{\mu}f$ is not a unit in $\mathcal{G}_{\mu}$, then $\mu=\operatorname{res}_{K}(v_{b,\delta})$ for some $b\in\operatorname{Z}(f)$, by Proposition 3.4. By Proposition 3.8, $\sigma(b)\in B(a,\delta)$ for some $\sigma\in\mathcal{D}$ and, obviously, $\sigma(b)\in\operatorname{Z}(f)$. Conversely, if $b\in\operatorname{Z}(f)\cap B(a,\delta)$, then $v_{b,\delta}=v_{a,\delta}$ and $\operatorname{in}_{\mu}f$ is not a unit in $\mathcal{G}_{\mu}$, by Proposition 3.4. ∎
In the Henselian case, this result was proved in [27, Proposition 3.16]
3.3. Proof of Theorem 3.3 when $\eta$ is valuation-transcendental
By Theorem 3.7, we may assume that $(K,v)$ is Henselian. Denote $\delta=\delta(\mu)$ and $\delta^{\prime}=\delta(\eta)$.
Take any $\phi\in\operatorname{KP}(\eta)$ and any $\chi\in\mathbf{t}(\mu,\eta)$. Since $\operatorname{in}_{\eta}\phi$ is a prime element, it is not a unit in $\mathcal{G}_{\eta}$. Hence, the two statements of Lemma 2.4 (ii) show that $\mu(\phi)<\eta(\phi)$ and $\operatorname{in}_{\mu}\chi\mid\operatorname{in}_{\mu}\phi$ in $\mathcal{G}_{\mu}$. Since $\operatorname{in}_{\mu}\chi$ is a prime element, $\operatorname{in}_{\mu}\phi$ is not a unit in $\mathcal{G}_{\mu}$ either. By Proposition 3.4, $\mu=\operatorname{res}_{K}(v_{a,\delta})$ and $\eta=\operatorname{res}_{K}(v_{b,\delta^{\prime}})$, for some $a,b\in\operatorname{Z}(\phi)$. Since $\phi$ is irreducible, we have $b=\sigma(a)$ for some $\sigma\in\mathcal{D}$.
The valuation $\overline{\eta}=v_{a,\delta^{\prime}}$ has the properties we are looking for. By Corollary 3.6, $\delta<\delta^{\prime}$, so that $\overline{\mu}<\overline{\eta}$. Finally, $\operatorname{res}_{K}(\overline{\eta})=\eta$ by Proposition 3.8. $\Box$
The following result is a consequence of Theorem 3.3 and Proposition 3.8.
Corollary 3.12.
Let $\mu,\eta\in\mathcal{V}$ be valuation-transcendental such that $\mu=\operatorname{res}_{K}(v_{a,\delta})$ and $\eta=\operatorname{res}_{K}(v_{b,\epsilon})$, for some pairs $(a,\delta),(b,\epsilon)\in\overline{K}\times\Lambda$. Then
$\mu\leq\eta\quad\Longleftrightarrow\quad\delta\leq\epsilon\ \mbox{ and }\ \sigma(b)\in B(a,\delta)\ \mbox{ for some }\sigma\in\mathcal{D}.$
3.4. Lifting increasing families of valuations
Definition 3.13.
Let $\mathcal{C}=\{\mu_{i}\}_{i\in I}$ be an increasing family of valuations in $\mathcal{V}$. We say that an increasing family of valuations $\{\overline{\mu}_{j}\}_{j\in J}$ in $\overline{\mathcal{V}}$ is an extension of $\mathcal{C}$ if
[…]
Theorem 3.14.
Let $\mathcal{C}=\{\mu_{i}\}_{i\in I}$ be an increasing family of valuations in $\mathcal{V}$, all of them valuation-transcendental. For a given $i_{0}\in I$, fix $\,\overline{\mu}_{i_{0}}\in\overline{\mathcal{V}}$ such that $\operatorname{res}_{K}(\overline{\mu}_{i_{0}})=\mu_{i_{0}}$. Then, there exists an increasing family of valuations $\,\overline{\mathcal{C}}=\{\overline{\mu}_{i}\}_{i\in I}$ in $\overline{\mathcal{V}}$ extending $\,\mathcal{C}$.
Proof.
By Theorem 3.7, we may assume that $(K,v)$ is Henselian. Consider the set $\mathcal{F}$ formed by pairs $\left(J,\{\overline{\mu}_{j}\}_{j\in J}\right)$ where $i_{0}\in J\subseteq I$ and $\{\overline{\mu}_{j}\}_{j\in J}$ is an increasing family of valuations extending $\{\mu_{j}\}_{j\in J}$. Define an ordering on $\mathcal{F}$ in the obvious way:
$\left(J,\{\overline{\mu}_{j}\}_{j\in J}\right)\leq\left(J^{\prime},\{\overline{\mu}^{\prime}_{j}\}_{j\in J^{\prime}}\right)\ \Longleftrightarrow\ J\subseteq J^{\prime}\ \mbox{ and }\ \overline{\mu}_{j}=\overline{\mu}^{\prime}_{j}\ \ \forall j\in J.$
We have ($\left(\{i_{0}\},\{\overline{\mu}_{i_{0}}\}\right)\in\mathcal{F}$, so $\mathcal{F}\neq\emptyset$. Moreover, it is easy to check that every totally ordered subset of $\mathcal{F}$ admits a majorant (taking the union over the sets $J$’s in the family). Hence, by Zorn’s Lemma, $\mathcal{F}$ admits maximal elements. Let us show that if $F_{\max}=\left(J,\{\overline{\mu}_{j}\}_{j\in J}\right)$ is maximal, then $I=J$. This will conclude the proof. For all $j\in J$, write $\delta_{j}:=\delta(\mu_{j})$ and $\overline{\mu}_{j}=v_{a_{j},\delta_{j}}$ for some $a_{j}\in\overline{K}$. Suppose that $J\neq I$ and take $i\in I\setminus J$. Suppose that $i<j$ for some $j\in J$. For each $k\in J$, $k\leq j$, we have $\overline{\mu}_{k}=v_{a_{j},\delta_{k}}$. Define $\overline{\mu}_{i}=v_{a_{j},\delta(\mu_{i})}$. Setting $J^{\prime}=J\cup\{i\}$ we have $\left(J^{\prime},\{\overline{\mu}_{j}\}_{j\in J^{\prime}}\right)\in\mathcal{F}$ and
$\left(J,\{\overline{\mu}_{j}\}_{j\in J}\right)<\left(J^{\prime},\{\overline{\mu}_{j}\}_{j\in J^{\prime}}\right)$
and this is a contradiction to the maximality of $F_{\max}$ in $\mathcal{F}$. Now, suppose that $J<i$. Consider any $\phi\in\operatorname{KP}(\mu_{i})$. Then, $\operatorname{in}_{\mu_{i}}\phi$ is not a unit in $\mathcal{G}_{\mu_{i}}$ and, as argued in the proof of Theorem 3.3, $\operatorname{in}_{\mu_{j}}\phi$ is not a unit in $\mathcal{G}_{\mu_{j}}$ either, for all $j\in J$. By Proposition 3.11, each ball $B(a_{j},\delta_{j})$ contains some root of $\phi$. Since $\phi$ has a finite number of roots and these balls form a descending chain, there is some $a\in\operatorname{Z}(\phi)$ belonging to the intersection of all these balls. The valuation $\overline{\mu}_{i}=v_{a,\delta(\mu_{i})}$ restricts to $\mu_{i}$ by Proposition 3.4, and clearly satisfies $\overline{\mu}_{i}>\overline{\mu}_{j}$ for all $j\in J$. This contradicts the maximality of $F_{\max}$. ∎
3.5. Proof of Theorem 3.3 when $\eta$ is valuation-algebraic
If $\eta$ is valuation-algebraic, then it is the stable limit $\eta=\lim_{i\in I}\mu_{i}$ of some increasing family $\{\mu_{i}\}_{i\in I}$ of inner nodes in $\mathcal{V}$. Since $\mathcal{V}$ is a tree and $\mu,\mu_{i}<\eta$, these valuations are comparable. Now, $\mu>\mu_{i}$ for all $i$ would imply $\eta=\lim_{i\in A}\mu_{i}\leq\mu$, against our assumption. Thus, there exists some $i_{0}\in A$ such that $\mu<\mu_{i}<\eta$ for all $i\geq i_{0}$.
By Theorem 3.3 for the valuation-transcendental case, there exists some $\overline{\mu}_{i_{0}}$ extending $\mu_{i_{0}}$ such that $\overline{\mu}<\overline{\mu}_{i_{0}}$. Therefore, we can assume that $\mu=\mu_{i_{0}}$.
By Theorem 3.14, there exists an increasing family $\{\overline{\mu}_{i}\}_{i\geq i_{0}}$ of valuations in $\overline{\mathcal{V}}$ extending the increasing family $\{\mu_{i}\}_{i\geq i_{0}}$. Let $\overline{\mu}_{i}=v_{a_{i},\delta_{i}}$ for all $i\geq i_{0}$. We claim that the corresponding ultrametric balls have empty intersection:
$\bigcap\nolimits_{i\geq i_{0}}B(a_{i},\delta_{i})=\emptyset.$
Indeed, suppose that $a\in\overline{K}$ belongs to $B(a_{i},\delta_{i})$ for all $i\geq i_{0}$. This means $v(a-a_{i})\geq\delta_{i}$, or equivalently, $\overline{\mu}_{i}(x-a)=\delta_{i}$. Hence,
$\overline{\mu}_{i}(x-a)=\delta_{i}<\delta_{j}=\overline{\mu}_{j}(x-a),\quad\mbox{ for all }j>i\geq i_{0}.$
Hence, $g:=p_{K}(a)$ satisfies $\mu_{i}(g)<\mu_{j}(g)$ for all $j>i\geq i_{0}$, contradicting the fact that $\left\{\mu_{i}\right\}_{i\in I}$ has a stable limit.
Therefore, $\{\overline{\mu}_{i}\}_{i\geq i_{0}}$ has a stable limit $\overline{\eta}=\lim_{i\geq i_{0}}\overline{\mu}_{i}\in\overline{\mathcal{V}}$ [27, Thm. 3.23]. By construction, $\overline{\mu}_{i}<\overline{\eta}$ for all $i\geq i_{0}$. On the other hand, for all $f\in K[x]$ there is a sufficiently large index $j$ such that
$\overline{\eta}(f)=\overline{\mu}_{j}(f)=\mu_{j}(f)=\eta(f),$
so that $\operatorname{res}_{K}(\overline{\eta})=\eta$. $\Box$
4. Abstract key polynomials
For some $\mu\in\mathcal{V}$, let us fix an extension $\overline{\mu}$ to $\overline{K}[x]$. Let us describe a certain chain $\mathcal{B}(\overline{\mu})$ of balls ordered by decreasing inclusion, determined by $\overline{\mu}$.
If $\mu$ is valuation-transcendental, then $\overline{\mu}=v_{a,\delta}$ for some pair $(a,\delta)\in\overline{K}\times\Lambda$. In this case, we denote
$\mathcal{B}=\mathcal{B}(\overline{\mu}):=\left\{B\in\mathbb{B}\mid B\supseteq B(a,\delta)\right\}.$
If $\mu$ is valuation-algebraic, then it is the stable limit of some increasing family $\mathcal{C}=\{\mu_{i}\}_{i\in I}$. As we saw in Section 3.4, $\overline{\mu}$ is the stable limit of some extension $\{\overline{\mu}_{i}\}_{i\in I}$ of $\mathcal{C}$ to $\overline{K}[x]$. For all $i\in I$, let $\overline{\mu}_{i}=v_{a_{i},\delta_{i}}$ and denote $B_{i}=B(a_{i},\delta_{i})$. In this case, we denote
$\mathcal{B}=\mathcal{B}(\overline{\mu}):=\left\{B\in\mathbb{B}\mid B\supseteq B_{i}\mbox{ for some }i\in I\right\}.$
4.1. Construction of all abstract key polynomials
The goal of this section is to express all abstract key polynomials for $\mu$ as minimal polynomials of certain centers of balls in $\mathbb{B}$.
To this end, the following easy observation will be useful.
Lemma 4.1.
Let $Q$ be an abstract key polynomial for $\mu\in\mathcal{V}$. Suppose that $\mu<\eta$ and $\mu(Q)=\eta(Q)$, for some $\eta\in\mathcal{V}$. Then, $Q$ is an abstract key polynomial for $\eta$.
Proof.
By Lemma 2.8, $Q\in\operatorname{KP}(\mu_{Q})$ is a key polynomial of minimal degree for $\mu_{Q}$. Thus, $\deg Q=\deg(\mu_{Q})\leq\deg(\mu)$. Hence, the key polynomials in $\mathbf{t}(\mu,\eta)$ have degree greater than or equal to $\deg Q$. By the definition of $\mathbf{t}(\mu,\eta)$, we have
$\mu(f)=\eta(f)\quad\mbox{ for all }f\in K[x]\ \mbox{such that}\ \deg f<\deg Q.$
In particular, $\mu\left(\partial_{s}Q\right)=\eta\left(\partial_{s}Q\right)$ and $\mu\left(\partial_{s}f\right)=\eta\left(\partial_{s}f\right)$, for all $s\in\mathbb{N}$. Therefore,
$\epsilon_{\eta}(f)=\epsilon_{\mu}(f)<\epsilon_{\mu}(Q)=\epsilon_{\eta}(Q).$
This proves that $Q$ is an abstract key polynomial for $\eta$. ∎
Theorem 4.2.
For every $\mu\in\mathcal{V}$ we have
$\Psi(\mu)=\bigcup_{B\in\mathcal{B}}p_{K}\left(\operatorname{Min}_{K}B\right).$
Proof.
Let us first assume that $\mu$ is valuation-transcendental.
For any $Q\in\Psi(\mu)$, let $\delta^{\prime}=\epsilon_{\mu}(Q)$ and take an optimizing root $c$ of $Q$; that is, $v_{b,\delta}(x-c)=\delta^{\prime}\leq\delta$, for some extension $v_{b,\delta}$ of $\mu$ to $\overline{K}[x]$.
By Proposition 3.8, we have $\sigma(b)\in B(a,\delta)$ for some $\sigma\in\mathcal{D}$. Hence,
$B:=B(\sigma(c),\delta^{\prime})\supseteq B(\sigma(b),\delta)=B(a,\delta),$
so that $B\in\mathcal{B}$. Now, since $\sigma(c)$ is an optimization root of $Q$ too, Theorem 2.10 shows that $\sigma(c)\in\operatorname{Min}_{K}B$. This proves that
(6) $\Psi(\mu)\subseteq\bigcup_{B\in\mathcal{B}}p_{K}\left(\operatorname{Min}_{K}B\right).$
For the converse, take any ball $B\in\mathbb{B}$ with $B=B(a^{\prime},\delta^{\prime})\supseteq B(a,\delta)$. For any $c\in\operatorname{Min}_{K}B$ we claim that $Q:=p_{K}(c)$ is an abstract key polynomial for $\mu$. Indeed, let us show that for every $f\in K[x]$, the condition $\deg f<\deg Q$ implies $\epsilon_{\mu}(f)<\epsilon_{\mu}(Q)$.
Since $a\in B$, we have $B=B(a,\delta^{\prime})$. Since $\delta^{\prime}\leq\delta$, for every $b\in\overline{K}$ we have
(7) $b\notin B\ \Longleftrightarrow\ v(b-a)<\delta^{\prime}\ \Longleftrightarrow\ v_{a,\delta}(x-b)<\delta^{\prime}.$
Since $c\in\operatorname{Min}_{K}B$, the condition $\deg f<\deg Q$ implies that $f$ has no root in $B$. In particular, for every root $b$ of $f$, we have $v_{a,\delta}(x-b)<\delta^{\prime}$, by (7). Since $\mu=\operatorname{res}_{K}(v_{a,\delta})$ and $v(a-c)\geq\delta^{\prime}$, Theorem 2.6 shows that
$\epsilon_{\mu}(f)<\delta^{\prime}\leq v_{a,\delta}(x-c)\leq\epsilon_{\mu}(Q)$
and this concludes the proof.
Let us now deal with the valuation-algebraic case.
Let $Q$ be an abstract key polynomial for $\mu$. Since $Q$ is an abstract key polynomial for $\mu_{Q}$, Lemma 4.1 shows that $Q$ is an abstract key polynomial for $\mu_{i}$ for a sufficiently large $i\in I$ for which $\mu_{Q}<\mu_{i}<\mu$.
By the valuation-transcendental case, $Q=p_{K}(c)$ for some $c\in\operatorname{Min}_{K}B$, for some ball $B$ containing $B_{i}$. Since $B\in\mathcal{B}$, this proves the inclusion (6).
For the converse, take any ball $B\in\mathbb{B}$ with $B\supseteq B_{i}$ for some $i\in I$, and take $Q=p_{K}(c)$ for some $c\in\operatorname{Min}_{K}B$. Take $j\in I$, $j>i$, large enough to stabilize $Q$; that is, $\mu_{j}(Q)=\mu(Q)$. By the valuation-transcendental case, $Q$ is an abstract key polynomial for $\mu_{j}$. By Lemma 4.1, $Q$ is an abstract key polynomial for $\mu$. ∎
Finally, let us recall [7, Theorem 5.3], which describes the balls associated to the truncated valuation $\mu_{Q}$, for every abstract key polynomial $Q$ for $\mu$.
Proposition 4.3.
For any $\mu\in\mathcal{V}$, take $Q=p_{K}(c)$ for some $c\in\operatorname{Min}_{K}B$, for some $B\in\mathcal{B}$. If $\overline{\mu}(x-c)=\epsilon_{\mu}(Q)$, then $\mu_{Q}=\operatorname{res}_{K}(v_{c,\epsilon_{\mu}(Q)})$.
4.2. Construction of a complete set of abstract key polynomials
Consider a set $\mathcal{S}\subseteq\Psi=\Psi(\mu)$ of abstract key polynomials for $\mu$. We say that $\mathcal{S}$ is complete if for every $f\in K[x]\setminus K$, there exists $Q\in\mathcal{S}$ such that
$\deg Q\leq\deg f\quad\mbox{and}\quad\mu_{Q}(f)=\mu(f).$
The existence of complete sets of abstract key polynomials was proved in [21], where it is described too, how to construct some sort of “minimal” complete sets.
For all $m\in\mathbb{N}$, let $\Psi_{m}\subseteq\Psi$ be the subset of polynomials of degree $m$. A complete set $\mathcal{S}$ can be constructed as follows. First, consider a pre-ordering on $\Psi$ according to the $\epsilon_{\mu}$-values:
$Q\leq Q^{\prime}\quad\Longleftrightarrow\quad\epsilon_{\mu}(Q)\leq\epsilon_{\mu}(Q^{\prime}).$
Then, inside each $\Psi_{m}\neq\emptyset$, take any totally ordered cofinal subset $\mathcal{S}_{m}\subseteq\Psi_{m}$. We then set $\mathcal{S}=\bigcup_{m}\mathcal{S}_{m}$.
The aim of this section is to reinterpret these results in terms of ultrametric balls. To this end, we introduce a relevant concept.
Lemma-Definition 4.4.
We say that $B=B(b,\gamma)\in\mathcal{B}$ is a $\mu$-optimal ball if it satisfies the following equivalent conditions
  1. (a)
    There exists $c\in\operatorname{Min}_{K}B$ such that  $\operatorname{res}_{K}(v_{b,\gamma})=\mu_{Q}$, where $Q=p_{K}(c)$.
  2. (b)
    There exists $c\in\operatorname{Min}_{K}B$ such that  $\epsilon_{\mu}(Q)=\gamma$, where $Q=p_{K}(c)$.
Proof.
Let us show that (a) $\Rightarrow$ (b). Since $\operatorname{res}_{K}(v_{c,\gamma})=\operatorname{res}_{K}(v_{b,\gamma})=\mu_{Q}$, we have $\gamma=\epsilon_{\mu_{Q}}(Q)$ by Theorem 2.6. On the other hand, $\epsilon_{\mu}(Q)=\epsilon_{\mu_{Q}}(Q)$.
Let us show that (b) $\Rightarrow$ (a). Let $\overline{\nu}=v_{b,\gamma}$. Since $B\in\mathcal{B}$, we have $\overline{\nu}\leq\overline{\mu}$. The condition $\epsilon_{\mu}(Q)=\gamma$ implies $\overline{\mu}(x-c)\leq\gamma$, by Theorem 2.6. Since $v(c-b)\geq\gamma$, we deduce that $\gamma=\overline{\nu}(x-c)\leq\overline{\mu}(x-c)$. Therefore, $\overline{\mu}(x-c)=\gamma$ and Proposition 4.3 shows that $\operatorname{res}_{K}(v_{c,\gamma})=\mu_{Q}$. Since $v_{c,\gamma}=v_{b,\gamma}$, we are done. ∎
Let us denote the subset of all $\mu$-optimal balls by
$\mathcal{B}^{\operatorname{opt}}=\mathcal{B}^{\operatorname{opt}}(\overline{\mu})\subseteq\mathcal{B}.$
For any $\mu$-optimal $B=B(b,\gamma)$ we define the set of optimal elements in $B$ as:
$\operatorname{Opt}(B):=\{c\in\operatorname{Min}_{K}B\mid\epsilon_{\mu}(Q)=\gamma,\mbox{ where }Q=p_{K}(c)\}.$
By Lemma-Definition 4.4, if $B,B^{\prime}\in\mathcal{B}^{\operatorname{opt}}$, we have
$p_{K}\left(\operatorname{Opt}(B)\right)\cap p_{K}\left(\operatorname{Opt}(B^{\prime})\right)\neq\emptyset\quad\Longrightarrow\quad B=B^{\prime},$
because the two balls have the same radius.
For $m\in\mathbb{N}$, we denote the subsets of those balls $B$ such that $\deg_{K}B=m$ by
$\mathcal{B}_{m}\subseteq\mathcal{B},\qquad\mathcal{B}^{\operatorname{opt}}_{m}\subseteq\mathcal{B}^{\operatorname{opt}}.$
Lemma 4.5.
Let $m\in\mathbb{N}$ such that $\mathcal{B}_{m}\neq\emptyset$. Then, $\mathcal{B}^{\operatorname{opt}}_{m}$ is a cofinal family in $\mathcal{B}_{m}$, with respect to descending inclusion.
Proof.
Suppose that $B=B(b,\epsilon)\in\mathcal{B}_{m}$ is not $\mu$-optimal. Take any $c\in\operatorname{Min}_{K}B$ and let $Q=p_{K}(c)$, $\gamma=\epsilon_{\mu}(Q)$. Clearly,
$\epsilon=v_{b,\epsilon}(x-c)\leq\overline{\mu}(x-c)\leq\gamma.$
Since $B$ in not $\mu$-optimal, we have necessarily $\epsilon<\gamma$. By Theorem 2.6, there exists some $c^{\prime}\in\operatorname{Z}(Q)$ such that $\overline{\mu}(x-c^{\prime})=\gamma$. Consider the ball $B^{\prime}:=B(c^{\prime},\gamma)$. The proof of the lemma will be complete if we show that $B^{\prime}$ satisfies:
(8) $B^{\prime}\in\mathcal{B}\quad\mbox{and}\quad B\supsetneq B^{\prime}.$
Indeed, $B\supsetneq B^{\prime}$ implies $\deg_{K}B\leq\deg_{K}B^{\prime}$. Since $\deg_{K}c^{\prime}=m$, we deduce that $\deg_{K}B^{\prime}=m$ and $c^{\prime}\in\operatorname{Min}_{K}B^{\prime}$. Thus, $B^{\prime}$ is $\mu$-optimal because $c^{\prime}$ satisfies the condition (b) of Lemma-Definition 4.4.
Let us first prove (8) when $\mu$ is valuation-transcendental. In this case, $\mathcal{B}$ is the set of closed balls containing $B(a,\delta)$, for some pair $(a,\delta)\in\overline{K}\times\Lambda$ such that $\overline{\mu}=v_{a,\delta}$. Since, $B\supseteq B(a,\delta)$, we have $B=B(a,\epsilon)$. Since $v_{a,\delta}(x-c^{\prime})=\gamma\leq\delta$, we see that
$v(a-c^{\prime})\geq\gamma>\epsilon.$
Hence, $c^{\prime}\in B$ and $a\in B^{\prime}$; or equivalently, $B\supsetneq B^{\prime}\supseteq B(a,\delta)$. This proves (8).
Now, suppose that $\mu$ is valuation-algebraic. Then, $\overline{\mu}$ is the stable limit of some increasing family $\left\{\overline{\mu}_{i}\right\}_{i\in I}$. For all $i\in I$, let $B_{i}=B(a_{i},\delta_{i})$, where $\overline{\mu}_{i}=v_{a_{i},\delta_{i}}$. We may take $i\in I$ large enough to have $B\supseteq B_{i}$ and $\epsilon_{\mu_{i}}(Q)=\epsilon_{\mu}(Q)$, simultaneously. Then, the above arguments show that $B\supsetneq B^{\prime}\supseteq B_{i}$. This proves (8) too. ∎
Corollary 4.6.
If $\mathcal{B}_{m}$ has a last (smallest) ball $B$, then $B$ is $\mu$-optimal.
As a consequence of Lemma 4.5, we may strengthen Theorem 4.2 as follows.
Theorem 4.7.
Let $\mu\in\mathcal{V}$. We can express $\Psi(\mu)$ as the disjoint union:
$\Psi(\mu)=\bigcup_{B\in\mathcal{B}^{\operatorname{opt}}}p_{K}\left(\operatorname{Opt}(B)\right).$
Now, it is easy to extract a minimal complete set of abstract key polynomials for $\mu$ as follows. Consider the sequence of all degrees of the balls in $\mathcal{B}$:
$1=m_{0}<m_{1}<\cdots<m_{n}<\cdots$
This sequence is finite if $\deg(\mu)$ is finite, the maximal degree being $\deg(\mu)$ in this case. The sequence is infinite only when $\mu$ is the stable limit of an increasing family $\{\mu_{i}\}_{i\in I}$ such that $\deg(\mu_{i})$ are unbounded.
We may split $\Psi(\mu)$ as the disjoint union $\bigcup_{n}\Psi_{m_{n}}$. Take for each $m_{n}$ a cofinal family of optimal balls of $\mathcal{B}_{m_{n}}$. For each optimal ball $B$ in this family, choose (only) one element $c\in\operatorname{Opt}(B)$ and consider the abstract key polynomial $p_{K}(c)$.
Let $\mathcal{S}_{m_{n}}\subseteq\Psi_{m_{n}}$ be the family of abstract key polynomials obtained in this way. Since the $\epsilon_{\mu}$-values of these abstract key polynomials are all different, the family $\mathcal{S}_{m_{n}}$ is totally ordered. Thus, $\mathcal{S}:=\bigcup_{n}\mathcal{S}_{m_{n}}$ is a complete set of abstract key polynomials for $\mu$.
5. Key polynomials
In this section, $\mu\in\mathcal{V}$ is assumed to be valuation-transcendental; or equivalently, $\operatorname{KP}(\mu)\neq\emptyset$ (Theorem 2.2). The first paper describing a relationship between balls and key polynomials for $\mu$ was [25]. As first examples of such a relationship let us quote the following facts.
Proposition 5.1.
If $\mu=\operatorname{res}_{K}(v_{a,\delta})$ for some $(a,\delta)\in\overline{K}\times\Lambda$, then
  • $\Psi(\mu)\cap{\rm KP}(\mu)=p_{K}\left(\operatorname{Min}_{K}B(a,\delta)\right)$.
  • $\operatorname{KP}(\mu)\subseteq p_{K}\left(B(a,\delta)\right)$.
Proof.
The first item is an immediate consequence of Theorem 2.10. The second item follows from Proposition 3.11 and the fact that key polynomials for $\mu$ are prime elements in the graded algebra $\mathcal{G}_{\mu}$. ∎
The next result is a useful way to characterize key polynomials for $\mu$. It will be the main tool used in this section.
Proposition 5.2.
The set of key polynomials for $\mu$ is given by
${\rm KP}(\mu)=\bigcup_{\mu<\eta}\mathbf{t}(\mu,\eta).$
Remark 5.3.
If we let the above union run through a set of representatives of the tangent directions of $\mu$, then this union is disjoint by Lemma 2.5. If $\mu=\operatorname{res}_{K}(v_{a,\delta})$ for some $(a,\delta)\in\overline{K}\times\Lambda$, we will show below that this disjoint union is in one-to-one correspondence with the partition of the ball $B(a,\delta)$ into open balls of radius $\delta$, up to the action of the decomposition group.
Proof of Proposition 5.2.
By Lemma 2.4, the set on the right hand side is contained in the set on the left hand side. Take any $\phi\in{\rm KP}(\mu)$ and any $\gamma\in\Gamma$ such that $\gamma>\mu(\phi)$. Consider the ordinary augmentation $\eta=[\mu;\,\phi,\gamma]$, defined on $\phi$-expansions as
(9) $\eta\left(f_{0}+f_{1}\phi+\ldots+f_{r}\phi^{r}\right):=\min_{0\leq i\leq r}\{\mu(f_{i})+i\gamma\}.$
Thus, $\mu(\phi)<\gamma=\eta(\phi)$ and $\mu(f)=\eta(f)$ for all $f\in K[x]$ with $\deg(f)<\deg(\phi)$. Consequently, $\phi\in\textbf{t}(\mu,\eta)$ and this concludes the proof. ∎
Lemma 5.4.
Take $\mu,\eta\in\mathcal{V}$ such that $\mu<\eta$. Let $\delta=\delta(\mu)$ and $\eta=\operatorname{res}_{K}(v_{b,\epsilon})$ for some pair $(b,\epsilon)\in\overline{K}\times\Lambda$. For all nonzero $f\in K[x]$ we have
$\mu(f)<\eta(f)\ \Longleftrightarrow\ f\mbox{ admits a root in }B^{\circ}(b,\delta).$
Proof.
By Theorem 3.1, there exists $a\in\overline{K}$ such that $\operatorname{res}_{K}(v_{a,\delta})=\mu$ and $v_{a,\delta}<v_{b,\epsilon}$. It is easy to see that for any element $c\in\overline{K}$ we have
$v_{a,\delta}(x-c)<v_{b,\epsilon}(x-c)\ \Longleftrightarrow\ c\in B^{\circ}(b,\delta).$
Hence the result follows. ∎
As a corollary of Lemma 5.4, we obtain the following result.
Corollary 5.5.
Assume that $\mu=\operatorname{res}_{K}(v_{a,\delta})$. For a nonzero $f\in K[x]$, there exists a valuation $\eta\in\mathcal{V}$ such that $\mu<\eta$ and $\mu(f)<\eta(f)$ if and only if $f$ admits a root in $B^{\circ}(b,\delta)$ for some $b\in B(a,\delta)$.
Proof.
Assume that $\mu(f)<\eta(f)$ for some $\mu<\eta$ in $\mathcal{V}$. We can assume, without loss of generality, that $\eta$ is valuation-transcendental. By Theorem 3.1, there exists an extension $v_{b,\epsilon}$ of $\eta$ such that $v_{a,\delta}<v_{b,\epsilon}$. By the previous lemma, $f$ admits a root in $B^{\circ}(b,\delta)$; on the other hand, $b\in B(a,\delta)$.
The converse follows from Lemma 5.4 by taking $\eta=\operatorname{res}_{K}(v_{b,v(c-b)})$, where $c$ is a root of $f$ in $B^{\circ}(b,\delta)$ for some $b\in B(a,\delta)$. ∎
The next result is a characterization of polynomials of ${\rm KP}(\mu)$ as minimal polynomials of suitable elements of $\mathbb{B}^{\circ}$.
Theorem 5.6.
Assume that $\mu=\operatorname{res}_{K}(v_{a,\delta})$ for some $(a,\delta)\in\overline{K}\times\Lambda$. Set
$\mathcal{U}=\{c\in B(a,\delta)\mid c\in\operatorname{Min}_{K}B^{\circ}(c,\delta)\}.$
Then, $\operatorname{KP}(\mu)=p_{K}(\mathcal{U})$.
Proof.
Take any valuation $\eta$ such that $\mu<\eta$. We can suppose that $\eta$ is valuation-transcendental. By Corollary 5.5, the elements $f\in K[x]$ for which $\mu(f)<\eta(f)$ are exactly those having a root in the ball $B^{\circ}(b,\delta)$ for some $b\in B(a,\delta)$.
Hence, every $\phi\in\mathbf{t}(\mu,\eta)$ has smallest degree among the polynomials having this property. This and Proposition 5.2 show that
$\operatorname{KP}(\mu)\subseteq p_{K}(\mathcal{U}).$
Now, take any $c\in\mathcal{U}$. In particular, $c\in B(a,\delta)$. Fix any $\delta^{\prime}>\delta$. Then, $\mu<\eta:=\operatorname{res}_{K}(v_{c,\delta^{\prime}})$. By Lemma 5.4, every nonzero $f\in K[x]$ such that $\mu(f)<\eta(f)$ has a root in $B^{\circ}(c,\delta)$. Hence, $\deg f\geq\deg_{K}c$. This shows that $p_{K}(c)\in\mathbf{t}(\mu,\eta)\subseteq\operatorname{KP}(\mu)$. Therefore, $p_{K}(\mathcal{U})\subseteq{\rm KP}(\mu)$ and this concludes the proof. ∎
As an immediate consequence of the above result we obtain the following.
Corollary 5.7.
Assume that $\mu=\operatorname{res}_{K}(v_{a,\delta})$ for some $(a,\delta)\in\overline{K}\times\Lambda$. Let
(10) $B(a,\delta)=\bigsqcup_{i\in I}B^{\circ}(a_{i},\delta)$
be the partition of $B(a,\delta)$ as a disjoint union of open balls of radius $\delta$. Then,
(11) ${\rm KP}(\mu)=\bigcup_{i\in I}{\rm KP}_{i}(\mu),\qquad\operatorname{KP}_{i}(\mu):=p_{K}\left(\operatorname{Min}_{K}B^{\circ}(a_{i},\delta)\right).$
We keep the notation of Corollary 5.7.
Lemma 5.8.
Take $\mu,\eta\in\mathcal{V}$ such that $\mu<\eta$. Assume that $\mu=\operatorname{res}_{K}(v_{a,\delta})$ for some $(a,\delta)\in\overline{K}\times\Lambda$ and $\eta=\operatorname{res}_{K}(v_{b,\epsilon})$ for some $b\in B(a,\delta)$ and $\epsilon>\delta$. Let $i\in I$ be uniquely determined by $B^{\circ}(b,\delta)=B^{\circ}(a_{i},\delta)$. Then
$\mathbf{t}(\mu,\eta)=\operatorname{KP}_{i}(\mu).$
Proof.
By definition, $f\in\mathbf{t}(\mu,\eta)$ if $\mu(f)<\nu(f)$ and $f$ has minimal degre with this property. By definition, $f\in{\rm KP}_{i}(\mu)$ if it has a root in $B^{\circ}(a_{i},\delta)$ and $f$ has minimal degree with this property. By Lemma 5.4, both conditions are equivalent. ∎
Therefore, in the decomposition (11), if ${\rm KP}_{i}(\mu)\cap{\rm KP}_{j}(\mu)\neq\emptyset$ for some $i,j\in I$, then ${\rm KP}_{i}(\mu)={\rm KP}_{j}(\mu)$. Let us see exactly which open balls can yield the same set $\operatorname{KP}_{i}(\mu)=\mathbf{t}(\mu,\eta)$. We recall that two key polynomials $f,g$ belong to the same tangent direction of $\mu$ if and only if $f\sim_{\mu}g$ (Lemma 2.4).
Lemma 5.9.
Let $f=p_{K}(c)$, $g=p_{K}(d)$ be the key polynomials for $\mu$ associated to $c,d\in\mathcal{U}$. Then, $f\sim_{\mu}g$ if and only if $B^{\circ}(c,\delta)=B^{\circ}(d,\delta)^{\sigma}$ for some $\sigma\in\mathcal{D}$.
Proof.
Suppose that $B^{\circ}(c,\delta)=B^{\circ}(d,\delta)^{\sigma}=B^{\circ}(\sigma(d),\delta)$ for some $\sigma\in\mathcal{D}$. Take $\epsilon:=v(c-\sigma(d))>\delta$ and $\eta:=\operatorname{res}(v_{c,\epsilon})$. By Lemma 5.8, $f,g\in\mathbf{t}(\mu,\eta)$, so that $f\sim_{\mu}g$ by Lemma 2.4. Conversely, suppose $f\sim_{\mu}g$ and take some valuation-transcendental $\eta$ on $K[x]$ such that $\mu<\eta$ and $f,g\in\mathbf{t}(\mu,\eta)$. By Theorem 3.1, there exist $b\in B(a,\delta)$ and $\epsilon\in\Gamma$, $\epsilon>\delta$, such that $\eta=\operatorname{res}(v_{b,\epsilon})$. By Lemma 5.4, $f$ and $g$ have a root in $B^{\circ}(b,\delta)$. Let $\sigma,\tau\in\mathcal{D}$ such that $\sigma(c),\tau(d)\in B^{\circ}(b,\delta)$. Then,
$B^{\circ}(c,\delta)^{\sigma}=B^{\circ}(b,\delta)=B^{\circ}(d,\delta)^{\tau}.$
This proves that $B^{\circ}(c,\delta)$ and $B^{\circ}(d,\delta)$ are conjugate under the action of $\mathcal{D}$. ∎
Therefore, the tangent directions of $\mu$ are in one-to-one correspondence with the decomposition (10), up to the action of $\mathcal{D}$. By Lemma 2.5, the tangent directions of $\mu$ parametrize the quotient set $\left(\mathcal{V}_{>\mu}\right)/\!\sim$, where $\sim$ is the equivalence relation:
$\nu\sim\eta\quad\Longleftrightarrow\quad(\mu,\nu]\cap(\mu,\eta]\neq\emptyset.$
For $\nu,\eta$ valuation-transcendental, this equivalence relation can be expressed as follows in terms of the decomposition (10). Let $\nu=\operatorname{res}(v_{b,\epsilon})$, $\eta=\operatorname{res}(v_{c,\epsilon^{\prime}})$ for some $b,c\in B(a,\delta)$ and $\epsilon,\epsilon^{\prime}>\delta$; then $\nu\sim\eta$ if and only if the balls $B^{\circ}(b,\delta)$ and $B^{\circ}(c,\delta)$ are conjugate under the action of $\mathcal{D}$.
The value-transcendental case
If $\mu$ is value-transcendental, then the graded algebra $\mathcal{G}_{\mu}$ has a unique homogeneous prime ideal and $\operatorname{KP}(\mu)/\!\sim_{\mu}$ is a one-element set [14, Theorem 4.2]. By Lemma 2.4, all tangent directions of $\mu$ coincide.
Therefore, the results of this section take a very simple form in this case. The extensions of $\mu$ to $\overline{K}[x]$ are of the form $v_{a,\delta}$ with $\delta\in\Lambda\setminus\Gamma$. Hence,
$B(a,\delta)=B^{\circ}(a,\delta).$
Thus, Theorem 5.6 and Corollary 5.7 are coherent with the fact that there is a unique tangent direction of $\mu$.
6. Limit key polynomials
In this Section, we consider the problem of characterizing limit key polynomials in terms of elements of $\mathbb{B}$.
Let $\mathcal{C}=\{\mu_{i}\}_{i\in I}$ be an increasing family of valuations in $\mathcal{V}$ admitting no last element. By Theorem 3.14, there exists an increasing family of valuations $\overline{\mathcal{C}}=\{\overline{\mu}_{i}\}_{i\in I}$ in $\overline{\mathcal{V}}$ extending $\mathcal{C}$. Thus, for each $i\in I$ we have $\overline{\mu}_{i}=v_{a_{i},\delta_{i}}$, for some pair $(a_{i},\delta_{i})\in\overline{K}\times\Lambda$. We obtain in this way a decreasing family of closed balls
$\{B_{i}\}_{i\in I,\qquad}B_{i}:=B(a_{i},\delta_{i})\in\mathbb{B}.$
It is well known [27, Proposition 2.12] that
$\operatorname{KP}_{\infty}(\mathcal{C})\neq\emptyset\quad\Longleftrightarrow\quad\bigcap_{i\in I}B_{i}\neq\emptyset.$
Let us reproduce a short proof of this result.
Proposition 6.1.
Let $B=\bigcap_{i\in I}B_{i}$. A polynomial $f\in K[x]$ is $\mathcal{C}$-unstable if and only if $\,\operatorname{Z}(f)\cap B\neq\emptyset$.
Proof.
Take $f\in K[x]$. By Lemma 2.4,(ii) we have
$f\ \mbox{ is $\,\mathcal{C}$-unstable}\quad\Longleftrightarrow\quad\operatorname{in}_{\mu_{i}}f\ \mbox{ is not a unit in $\mathcal{G}_{\mu_{i}}$ for all }i\in I.$
Now, Proposition 3.11 shows that the right hand side condition is equivalent to $\,\operatorname{Z}(f)\cap B_{i}\neq\emptyset$ for all $i\in I$. Since $f$ has a finite number of zeros, this is equivalent to $\,\operatorname{Z}(f)\cap B\neq\emptyset\,$ too. ∎
As a consequence, we obtain a result for limit key polynomials which is coherent with Theorem 4.2 for abstract key polynomials and Theorem 5.6 for key polynomials.
Theorem 6.2.
Let $B=\bigcap_{i\in I}B_{i}$. Then, $\operatorname{KP}_{\infty}(\mathcal{C})=p_{K}(\operatorname{Min}_{K}B)$.
Proof.
We need only to consider objects of minimal degree satisfying the equivalent conditions of Proposition 6.1. ∎
Remark 6.3.
Given an increasing family $\mathcal{C}=\{\mu_{i}=(v_{a_{i},\delta_{i}})_{|K[x]}\}$ of valuations in $\mathcal{V}$ as above such that the index set $I$ is well-ordered, if we choose $a_{i}\in B_{i}$ so that $v(a_{i}-a_{j})=\delta_{i}$ for $i<j$, then $E=\{a_{i}\}_{i\in I}$ is a pseudo-convergent sequence in $\overline{K}$ (see [22] and also [6, p. 283] for this approach) and $B$ is precisely the set of pseudo-limits of $E$ in $\overline{K}$. The sequence $E$ is either of algebraic type or of transcendental type in the sense of Kaplansky ([11]), if either $B$ is non-empty or not, respectively ([23, 24]). If $B\not=\emptyset$, then, by Proposition 6.1, the $\mathcal{C}$-unstable polynomials $f\in K[x]$ are those having a root which is a pseudo-limit of $E$ and, by Theorem 6.2, the limit key polynomials for $\mathcal{C}$ are the minimal polynomials of some pseudo-limit of $E$ having minimal degree. If $B=\emptyset$ and thus every $f\in K[x]$ is $\mathcal{C}$-stable, by Corollary 3.5 for each such an $f$ we have $\mu_{i}(f)=v(f(a_{i}))$ for all $i\in I$ sufficiently large and therefore $\mu=\lim(\mathcal{C})$ as defined in (3) is precisely equal to the valuation $v_{E}$ associated to $E$, defined as $v_{E}(f)=v(f(a_{i}))$ for all $i$ sufficiently large (see [23, 24] for other relevant results regarding the valuation $v_{E}$).
7. On Mac Lane–Vaquié chains
A Mac Lane-Vaquié chain (abbreviated MLV chain) of a given $\mu\in\mathcal{V}$ is a finite, or countably infinite, totally ordered sequence in $\mathcal{V}$ upper bounded by $\mu$:
$\mu_{0}\,<\,\mu_{1}\,<\,\cdots\,<\,\mu_{n}\,<\,\cdots\leq\mu$
satisfying the following conditions:
(MLV1)  $1=\deg(\mu_{0})\,<\,\deg(\mu_{1})\,<\,\cdots\,<\,\deg(\mu_{n})\,<\,\cdots$.
(MLV2)  Each step $\mu_{n-1}<\mu_{n}$ is an augmentation of one of the following kinds:
$\bullet$  ordinary augmentation: $\mu_{n}=[\mu_{n-1};\,\phi_{n},\gamma_{n}]$, for some $\phi_{n}\in\operatorname{KP}(\mu_{n-1})$ and some $\gamma_{n}\in\Lambda$, $\gamma_{n}>\mu_{n-1}(\phi_{n})$ (see the definition in (9)).
$\bullet$  limit augmentation: there is a well-ordered increasing family $\mathcal{C}_{n-1}=\{\rho_{i}\}_{i\in I_{n-1}}$ of valuations of constant degree, such that $\mu_{n-1}=\min\left(\mathcal{C}_{n-1}\right)$ and $\mu_{n}$ is the limit augmentation:  $\mu_{n}=[\mathcal{C}_{n-1};\,\phi_{n},\gamma_{n}]$, for some $\phi_{n}\in\operatorname{KP}_{\infty}(\mathcal{C}_{n-1})$ and some $\gamma_{n}\in\Lambda$, $\gamma_{n}>\rho_{i}(\phi_{n})$ for all $i\in I_{n-1}$.
(MLV3)  $\phi_{n}\not\in\mathbf{t}(\mu_{n},\mu)$ for all $\mu_{n}<\mu$.
(MLV4)  If $\mu$ is valuation-transcendental, then the sequence has $\mu$ as its last valuation:
(12) $\mu_{0}\,<\,\mu_{1}\,<\,\cdots\,<\,\mu_{r}=\mu.$
If $\mu$ is valuation-algebraic of finite degree, then the sequence has a last valuation (say) $\mu_{r-1}$ and there is a well-ordered increasing family $\mathcal{C}_{r-1}$ of valuations of constant degree, such that $\mu_{r-1}=\min\left(\mathcal{C}_{r-1}\right)$ and $\mu$ is the stable limit of $\mathcal{C}_{r-1}$:
(13) $\mu_{0}\,<\,\mu_{1}\,<\,\cdots\,<\,\mu_{r-1}\,<\,\mu=\lim\left(\,\mathcal{C}_{r-1}\right).$
If $\mu$ is valuation-algebraic of infinite degree, then the sequence is infinite and $\mu=\lim_{n\in\mathbb{N}}\mu_{n}$ is the stable limit of the sequence.
The existence of such chains was proved by Mac Lane in the discrete rank-one case, and by Vaquié in the general case [13, 26]. Their relevance lies in the fact that they contain intrinsic information about $\mu$. For instance, in [15, Section 4] it is proved that the following data are common to all MLV chains of the same $\mu$:
$\bullet$  The length of the chain, called the depth of $\mu$. A valuation $\mu$ as in (12) or (13) has depth $r$, while a valuation-algebraic $\mu$ of infinite degree has infinite depth.
$\bullet$  The sequence of degrees  $1=\deg(\mu_{0})\,<\,\deg(\mu_{1})\,<\,\cdots\,<\,\deg(\mu_{n})\,<\,\cdots$.
$\bullet$  The type, ordinary or limit, of each augmentation $\mu_{n-1}<\mu_{n}$.
$\bullet$  The valuations $\mu_{n}$ such that $\mu_{n}<\mu_{n+1}$ is an ordinary augmentation.
The conditions (MLV1), (MLV2) and (MLV4) are sufficient to derive these “unicity” results. The condition (MLV3) is equivalent to
$\mu_{n}(\phi_{n})=\gamma_{n}=\mu(\phi_{n})\quad\mbox{for all }\ n.$
This implies that all nodes of the chain, except for (eventually) $\mu$, are residue-transcendental.
The goal of this section is to show that a MLV chain for $\mu$ can be read directly in the descending family $\mathcal{B}^{\operatorname{opt}}=\mathcal{B}^{\operatorname{opt}}(\overline{\mu})$ of $\mu$-optimal ultrametric balls, associated to any extension $\overline{\mu}$ of $\mu$ to $\overline{K}[x]$ in Section 4.2.
We need a well-known auxiliary observation. For the ease of the reader we give a short proof of it.
Lemma 7.1.
Let $\nu,\mu$ be valuations on $K[x]$ such that $\nu<\mu$. Let $\phi\in\operatorname{KP}(\mu)$ be a key polynomial for $\mu$ of minimal degree.
  1. (i)
    If $\phi\in\mathbf{t}(\nu,\mu)$, then $\mu=[\nu;\,\phi,\mu(\phi)]$.
  2. (ii)
    Suppose that $\nu$ is the initial valuation of an increasing family $\mathcal{C}=\left\{\rho_{i}\right\}_{i\in I}$ such that $\rho_{i}<\mu$ for all $i\in I$. If $\phi\in\operatorname{KP}_{\infty}(\mathcal{C})$, then $\mu=[\mathcal{C};\,\phi,\mu(\phi)]$.
Proof.
In case (i), we may consider the ordinary augmentation $\mu^{\prime}=[\nu;\,\phi,\mu(\phi)]$, because $\phi\in\mathbf{t}(\nu,\mu)\subseteq\operatorname{KP}(\nu)$ and $\nu(\phi)<\mu(\phi)$. In case (ii), we may consider the limit augmentation $\mu^{\prime}=[\mathcal{C};\,\phi,\mu(\phi)]$, because $\phi\in\operatorname{KP}_{\infty}(\mathcal{C})$ and $\rho_{i}(\phi)<\mu(\phi)$ for all $i\in I$. In both cases, $\mu^{\prime}=\mu$ because both valuations coincide on $\phi$-expansions. ∎
Let us recall the definition of $\mathcal{B}^{\operatorname{opt}}$. If $\mu$ is valuation-transcendental and $\overline{\mu}=v_{a,\delta}$, then the balls $B\in\mathcal{B}^{\operatorname{opt}}$ are $\mu$-optimal and satisfy $B\supseteq B(a,\delta)$. On the other hand, if $\mu$ is valuation-algebraic and $\overline{\mu}=\lim_{i\in I}v_{a_{i},\delta_{i}}$, then the balls $B\in\mathcal{B}^{\operatorname{opt}}$ are $\mu$-optimal and satisfy $B\supseteq B(a_{i},\delta_{i})$ for some $i\in I$.
Notation. For each $B=B(b,\epsilon)\in\mathcal{B}^{\operatorname{opt}}$, denote $\mu_{B}=\operatorname{res}_{K}(v_{b,\epsilon})$.
Consider the sequence of degrees over $K$ of all balls in $\mathcal{B}^{\operatorname{opt}}$:
(14) $1=m_{0}\,<\,m_{1}\,<\,\cdots\,<\,m_{n}\,<\,\cdots$
Denote by $\mathcal{B}^{\operatorname{opt}}_{m_{n}}\subseteq\mathcal{B}^{\operatorname{opt}}$ the subset of balls of degree $m_{n}$ over $K$.
Theorem 7.2.
For each $n\in\mathbb{N}_{0}$ with $\mathcal{B}^{\operatorname{opt}}_{m_{n}}\neq\emptyset$, take a well-ordered cofinal family $\mathcal{B}_{m_{n}}\subseteq\mathcal{B}^{\operatorname{opt}}_{m_{n}}$ with respect to descending inclusion, such that:
$\mathcal{B}_{m_{n}}=\{B\}\quad\mbox{whenever $\,\mathcal{B}^{\operatorname{opt}}_{m_{n}}$ has a last ball }B.$
Take $\mu_{n}:=\mu_{B_{n}}$, for each $B_{n}:=\min\left(\mathcal{B}_{m_{n}}\right)$. Then, the chain
$\mu_{0}\,<\,\mu_{1}\,<\,\cdots\,<\,\mu_{n}\,<\,\cdots$
is a Mac Lane-Vaquié chain of $\mu$.
Proof.
By Theorem 4.7 and Lemma-Definition 4.4, for each ball $B_{n}$ there exists $c\in\operatorname{Min}_{K}B_{n}$ such that $\phi_{n}:=p_{K}(c)$ is an abstract key polynomial for $\mu$ satisfying $\mu_{n}=\mu_{\phi_{n}}$. By Lemma 2.8, $\phi_{n}$ is a key polynomial of minimal degree of $\mu_{n}$, so that $m_{n}=\deg_{K}B_{n}=\deg\phi_{n}=\deg(\mu_{n})$. Thus, the chain satisfies (MLV1) because the sequence of degrees of the valuations $\mu_{n}$ coincides with the sequence (14).
Consider any $\mu_{n-1}<\mu_{n}$ such that $\mathcal{B}_{m_{n-1}}=\{B\}$. Since $B\supsetneq B_{n}$, Theorem 5.6 and Lemma 5.8 show that $\phi_{n}\in\mathbf{t}(\mu_{n-1},\mu_{n})$. By Lemma 7.1, $\mu_{n}$ is the ordinary augmentation $[\mu_{n-1};\,\phi_{n},\gamma_{n}]$, where $\gamma_{n}=\mu_{\phi_{n}}(\phi_{n})=\mu(\phi_{n})$.
Consider any $\mu_{n-1}<\mu_{n}$ such that $\mathcal{B}_{m_{n-1}}=\left\{B_{i}\right\}_{i\in I_{n-1}}$ is infinite. Then, $\mathcal{C}_{n-1}:=\left\{\mu_{B_{i}}\right\}_{i\in I_{n-1}}$ is an increasing family of valuations of constant degree, with $\mu_{n-1}=\min\left(\mathcal{C}_{n-1}\right)$. Since $B_{i}\supsetneq B_{n}$ for all $i\in I_{n-1}$, Theorem 6.2 shows that $\phi_{n}\in\operatorname{KP}_{\infty}(\mathcal{C}_{n-1})$. By Lemma 7.1, $\mu_{n}=[\mathcal{C}_{n-1};\,\phi_{n},\gamma_{n}]$, for $\gamma_{n}=\mu_{\phi_{n}}(\phi_{n})=\mu(\phi_{n})$.
This proves that the chain satisfies (MLV2) and (MLV3).
If $\mu$ is valuation-transcendental and $\overline{\mu}=v_{a,\delta}$, then the last ball $B(a,\delta)$ of $\,\mathcal{B}(\overline{\mu})$ is $\mu$-optimal (Corollary 4.6). Hence, if $\deg(\mu)=m_{r}$, we have $\mathcal{B}_{m_{r}}=\{B(a,\delta)\}$ and $\mu_{r}=\mu_{\phi_{r}}=\operatorname{res}_{K}(v_{a,\delta})=\mu$, by Lemma-Definition 4.4.
Finally, suppose that $\mu_{n-1}<\mu$ and $\mathcal{B}_{m_{n-1}}=\left\{B_{i}\right\}_{i\in I}$ satisfies $\bigcap_{i\in I}B_{i}=\emptyset$. Let $\mathcal{C}_{n-1}:=\left\{\mu_{B_{i}}\right\}_{i\in I}$. By Proposition 6.1, $\mathcal{C}_{n-1}$ has a stable limit. Since $\nu:=\lim\left(\,\mathcal{C}_{n-1}\right)$ is valuation-algebraic and $\nu\leq\mu$, we have $\nu=\mu$ by Theorem 2.2. ∎
Remark. In the Henselian case, Vaquié showed as well, how to derive an MLV chain of a valuation-transcendental $\mu=\operatorname{res}_{K}(v_{a,\delta})$, directly from the balls containing $B(a,\delta)$ [27, Section 4].